# Action Controlled Paraphrasing

### v1

@@ Abstract
Recent studies have demonstrated the potential to control paraphrase generation, such as through syntax, which has broad applications in various downstream tasks. However, these methods often require detailed parse trees or syntactic exemplars, which are not user-friendly. Furthermore, an inference gap exists, as control specifications are only available during training but not inference. In this work, we propose a new setup for controlled paraphrasing. Specifically, we represent user-intended actions as action tokens, allowing embedding and concatenating them with text embeddings, thus flowing together to a self-attention encoder for representation fusion. To address the inference gap, we introduce an optional action token as a placeholder that encourages the model to determine the appropriate action when control specifications are inaccessible. Experimental results show that our method successfully enables specific action-controlled paraphrasing and preserves the same or even better performance compared to conventional uncontrolled methods when actions are not given. Our findings thus promote the concept of optional action control for a more user-centered design via representation learning.
@@
Ning Shi1  Zijun Wu1  Lili Mou1,2 1Dept. Computing Science, Alberta Machine Intelligence Institute (Amii), University of Alberta 2Canada CIFAR AI Chair, Amii ning.shi@ualberta.ca  zijun4@ualberta.ca  doublepower.mou@gmail.com
@@ 1 Introduction
By incorporating attention mechanisms (Vaswani et al., 2017) and leveraging pre-trained language models (Devlin et al., 2019), the sequence-to-sequence (seq2seq) framework (Sutskever et al., 2014) has achieved impressive results (Prakash et al., 2016; Wang et al., 2019; Witteveen and Andrews, 2019; Niu et al., 2021). Some have explored other methodologies, including variational autoencoders (Kingma et al., 2014; Gupta et al., 2018; Bao et al., 2019; Roy and Grangier, 2019), generative adversarial networks (Goodfellow et al., 2014; Yang et al., 2019; Vizcarra and Ochoa-Luna, 2020), reinforcement learning (Li et al., 2018; Siddique et al., 2020), and information retrieval-based approaches (Kazemnejad et al., 2020).
Recent studies have gradually shifted to incorporate user control (Iyyer et al., 2018; Chen et al., 2019; Goyal and Durrett, 2020; Kumar et al., 2020; Hosking et al., 2022). These works aim to generate paraphrases that satisfy specific control requirements such as a syntax template (Iyyer et al., 2018) or a sentential exemplar (Chen et al., 2019). However, these methods often lack user-friendliness. They require users to provide detailed control specifications, even when such specifications are unnecessary. Additionally, an inference gap exists. It is often assumed that control specifications are always available, which is typically true only during training but not during inference or real-world application. These limitations highlight the need for more user-friendly and flexible approaches to controlled paraphrase generation.
In this work, we propose action-controlled paraphrasing as a new setup to address these challenges. Specifically, we represent users’ intentions as action tokens, indicating words to keep or paraphrase. They serve as additional supervision signals, can be automatically derived from the differences between source texts and target paraphrases without external toolkits or manual labor. We introduce action embeddings alongside word and positional embeddings, concatenating all three and feeding them into a self-attention encoder for representation fusion. To close the inference gap, we employ a special action token for the model to independently determine the appropriate actions without explicit guidance from the user. Experimental results demonstrate the effectiveness of our methods. On one hand, the specific action control allows users to impose constraints and promote diverse paraphrasing. On the other hand, the optional action control balances the trade-off between performance and constraints. Our approach also holds potential for applications such as lexical substitution (Qiang et al., 2023) and adversarial example generation (Iyyer et al., 2018), which targets the paraphrasing of specific fragments of text.
@@ 6 Conclusion
Overall, our work revolutionizes paraphrase generation by introducing action control, allowing users to selectively determine which words to keep or paraphrase. We propose an end-to-end procedure for generating actions and seamlessly incorporating them into the input encoding through representation learning. During training, the action generation process is fully automated, requiring no domain knowledge, external toolkits, or manual labor. The training strategy incorporates an optional action, enabling models to make independent decisions without user guidance, bridging the gap between training and inference. Empirical evaluations demonstrate the effectiveness of our approach under specific and optional action control. Specific action control achieves precise paraphrasing aligned with targets, and optional action control balances the trade-off between performance and constraints. Our framework offers flexibility for future approaches to adopt optional constraints, promoting user-centered designs and thus enhancing user experiences in paraphrase generation.
@@ Acknowledgment
We gratefully appreciate Bradley Hauer and Grzegorz Kondrak for sharing their pearls of wisdom. The research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), a Mitacs Accelerate project, the Amii Fellow Program, the Canada CIFAR AI Chair Program, an Alberta Innovates Program, and the Digital Research Alliance of Canada (alliancecan.ca).

### v2

@@ Abstract
Recent studies have demonstrated the potential to control paraphrase generation, such as through syntax, which has broad applications in various downstream tasks. However, these methods often require detailed parse trees or syntactic exemplars, countering human-like paraphrasing behavior in language use. Furthermore, an inference gap exists, as control specifications are only available during training but not during inference. In this work, we propose a new setup for controlled paraphrase generation. Specifically, we represent user intent as action tokens, embedding and concatenating them with text embeddings, thus flowing together into a self-attention encoder for representation fusion. To address the inference gap, we introduce an optional action token as a placeholder that encourages the model to determine the appropriate action independently when users’ intended actions are not provided. Experimental results show that our method successfully enables precise action-controlled paraphrasing and preserves or even enhances performance compared to conventional uncontrolled methods when actions are not given. Our findings promote the concept of action-controlled paraphrasing for a more user-centered design.
@@
Ning Shi  Zijun Wu Alberta Machine Intelligence Institute (Amii) Department of Computing Science University of Alberta, Edmonton, Canada {ning.shi,zijun4}@ualberta.ca
@@ 1 Introduction
Advanced neural approaches have significantly contributed to this field (Vaswani et al., 2017; Devlin et al., 2019), where the sequence-to-sequence framework (Sutskever et al., 2014) has achieved impressive results (Prakash et al., 2016; Wang et al., 2019; Witteveen and Andrews, 2019; Niu et al., 2021). Some have explored other methodologies, including variational autoencoders (Kingma et al., 2014; Gupta et al., 2018; Bao et al., 2019; Roy and Grangier, 2019), generative adversarial networks (Goodfellow et al., 2014; Yang et al., 2019; Vizcarra and Ochoa-Luna, 2020), reinforcement learning (Li et al., 2018; Siddique et al., 2020), and information retrieval (Kazemnejad et al., 2020).
Recent studies have gradually shifted to incorporate user control (Iyyer et al., 2018; Chen et al., 2019; Goyal and Durrett, 2020; Kumar et al., 2020; Hosking et al., 2022). These works aim to generate paraphrases that satisfy specific control requirements such as a syntax template (Iyyer et al., 2018) or a sentential exemplar (Chen et al., 2019). However, these methods often lack user-friendliness. They require users to provide detailed control specifications, even when such specifications are unnecessary. They are counter-intuitive to natural human behavior in paraphrase generation. Additionally, an inference gap exists. It is often assumed that control specifications are always available, which is typically true only during training but not inference or real-world application. These limitations highlight the need for more user-friendly and flexible approaches to controlled paraphrase generation.
In this work, we propose action-controlled paraphrasing as a new setup to address these challenges. Specifically, we represent users’ intentions as action tokens, indicating words to keep or paraphrase. As additional supervision signals, they can be automatically derived from the differences between source texts and target paraphrases without external toolkits or manual labor. We introduce action embeddings alongside word and positional embeddings, concatenating all three and feeding them into a self-attention encoder for representation fusion. To close the inference gap, we employ a special action token for the model to independently determine the appropriate actions without explicit guidance from the user. Experimental results demonstrate the effectiveness of our methods. On one hand, the specific action control allows users to impose constraints and promote diverse paraphrasing. On the other hand, the optional action control balances the trade-off between the control and performance. Our approach also holds potential for applications such as lexical substitution (Qiang et al., 2023) and adversarial example generation (Iyyer et al., 2018), which targets the paraphrasing of specific fragments of text.
@@ 6 Conclusion
Overall, we have presented action-controlled paraphrasing, allowing users to selectively determine which words to keep or paraphrase. We propose an end-to-end procedure for representing user intent as actions and seamlessly incorporating them into the input encoding through representation fusion. During training, the action generation process is fully automated, requiring no domain knowledge, external toolkits, or manual labor. To mirror human-like adaptability, we incorporate a special action token into the training strategy, enabling models to make independent decisions without user guidance and thus bridging the gap between training and inference. Empirical evaluations demonstrate the effectiveness of our approach under both specific and optional action control. Specific action control achieves precise paraphrasing aligned with input actions, and optional action control balances the trade-off between the control and performance. Our framework promotes the adoption of optional control in future research, aligning more closely with human cognitive processes, and enhancing user experiences in paraphrase generation.
@@ Acknowledgment
We gratefully appreciate Bradley Hauer, Grzegorz Kondrak, and Lili Mou for sharing their pearls of wisdom. This work is based on a course project for CMPUT651: Deep Learning for NLP at the University of Alberta.